OpenStudy (anonymous):

A sports car moving at constant speed travels 140m in 5.0s, then brakes and comes to a stop in 4.1s. What is the magnitude of its acceleration in m/s^2? What is the magnitude of its acceleration in g's?

OpenStudy (shane_b):

Acceleration is equal to the rate of change in velocity so first you need to calculate the car's initial velocity:\[\frac{140m}{5.0s}=28m/s\] Now you know the car is moving along at a constant 28 m/s (no acceleration at this point). It then brakes to a stop in 4.1s so calculate it's acceleration during that period:\[a=\frac{V_f-V_i}{\Delta t}=\frac{0m/s-28/ms}{4.1s}=-6.8m/s^2\]Convert that to g's to get the second part of the answer:\[\frac{-6.8m/s^2}{-9.8m/s^2}=0.69g\]